Let.

1             0             0

0             3             6

0             -1           -2


Verify that C 2 = C holds. Find the eigenvalues and eigenvectors of C. 


1
Expert's answer
2021-12-16T09:52:57-0500

"C^2=\begin{pmatrix}
 1 & 0&0 \\
 0 & 3&6\\
0&-1&-2
\end{pmatrix}\begin{pmatrix}
 1 & 0&0 \\
 0 & 3&6\\
0&-1&-2
\end{pmatrix}=\begin{pmatrix}
 1 & 0&0 \\
 0 & 10&22\\
0&-2&-2
\end{pmatrix}"


"\begin{vmatrix}
 1-\lambda & 0&0 \\
 0&3- \lambda& 6\\
0&-1&-2-\lambda
\end{vmatrix}=0"


"(1-\lambda)((3-\lambda)(-2-\lambda)+6)=0"

"\lambda_1=1"

"\lambda^2-\lambda=0"

"\lambda_2=0,\lambda_3=1"


for "\lambda=1" :

"2y+6z=0"

"-y-3z=0"

"y=-3z"


"u_1=\begin{pmatrix}
 1 \\
 -1\\
3
\end{pmatrix},u_2=\begin{pmatrix}
 1 \\
 1\\
-3
\end{pmatrix}"


for "\lambda=0" :

"x=0"

"3y+6z=0"

"-y-2z=0"

"y=-2z"


"u_3=\begin{pmatrix}
 0 \\
 -1\\
2
\end{pmatrix}"